OpenStudy (anonymous):

Find the volume of the regular pyramid. Square base with side 30.6ft; slant height 46.8 ft.

OpenStudy (anonymous):

first find the height of the pyramid using pythagoras theorem with 46.8 = hypotenuse and base = half of 30.6 = 15.3 then vol = (1/3) * area base * height

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

any more questions - just ask

OpenStudy (anonymous):

will do... just working on it... thanks .. when I get my answer will you verify it?

OpenStudy (anonymous):

is the height 44.23?

OpenStudy (anonymous):

hey that was quick

OpenStudy (amistre64):

height = h^2 + (30.6/2)^2 = (46.8)^2

OpenStudy (amistre64):

its a short drive lol

OpenStudy (amistre64):

h = 44.228

OpenStudy (amistre64):

volume = 30.3^2 * 44.228/3 = 13,535.0948 or so

OpenStudy (anonymous):

ok I did one right huh

OpenStudy (amistre64):

ugh...30.6 lol

OpenStudy (amistre64):

13,804.4434

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I'm frying your brain up!

OpenStudy (amistre64):

yes you did ;)

OpenStudy (anonymous):

ok final product 13,804.443 ft cubed

OpenStudy (amistre64):

yes, cubed

OpenStudy (amistre64):

like qbert, but different

OpenStudy (anonymous):

LOL